# MPMA: Preference Manipulation Attack Against Model Context Protocol

### v1

@@ 5 Experiment Setup
Attack Success Rate (ASR). This metric evaluates the attack effectiveness of the MPMA. It can be defined in EquationΒ 3:
@@
True Positive Rate (TPR). This metric measures the stealthiness of the MCP server by measuring the proportion of tool descriptions that are flagged as suspicious by either LLMs or human annotators. To conduct this assessment, we utilize the LLM-as-a-judge paradigm[41] and human evaluation to evaluate whether a given tool description may raise suspicion. For LLM-as-a-judge, five of the most widely used LLMs are employed for this assessment: Deepseek-V3 (Deepseek)[46], Claude-3.7-Sonnet (Claude)[5], GPT-4o[43], Grok-3 (Grok)[15], and Gemini-2.5 (Gemini)[53]. To enhance the domain knowledge of LLMs in assessing whether a tool description appears suspicious, this study adopts the widely used few-shot In-Context Learning (ICL)[39; 36] strategy. Specifically, we manually designed manipulative examples to serve as demonstrations to calibrate the LLMsβ suspicion detection. For human evaluation, three independent annotators are recruited. They label the tool description based on the instruction shown in AppendixΒ B. The calculation of the TPR metric can be defined in the following EquationΒ 4:
@@
where denotes the set of judge models, denotes the MCP server list, represents the judge prompt for stealthiness judging, and is the judging demonstration. The judging instructions and demonstrations can be seen in AppendixΒ B. Note that for human evaluation, the equation of should be where the means judge by human. A lower TPR indicates the attack can evade the censor of the LLM and human, which turns out to be more stealthy.
@@
Model. We evaluate the MPMA utilizing five widely adopted base LLMs for LLM agent: Deepseek-V3 (Deepseek)[46], Claude-3.7-Sonnet (Claude)[5], Gemini-2.5-flash (Gemini)[53], Qwen3-235B-A22B (Qwen3)[56], and GPT-4o[43].
@@
MCP Server. 8 commonly used MCP servers are employed in the experiments. These servers provide the following functionalities: weather information (Weather)[31], time information (Time)[29], MCP server installation assistance (Installer)[17], daily hot news (Hotnews)[16], web page content fetching (Fetch)[13], web-to-markdown conversion (Markdown)[32], cryptocurrency analysis (Crypto)[8], and web search (Search)[28]. The demonstration of tool description can be seen in AppendixΒ B.
@@
Dataset. For each MCP server, ten common queries corresponding to the MCP server are constructed for evaluation. More details can be seen in AppendixΒ B.
@@
Implementation Details. To simulate a competitive environment, five additional competing MCP servers with the same functionality are included alongside the malicious MCP server. These competing servers share the same name, and their descriptions are paraphrased using GPT-4o[43] to ensure diversity. In the main experiments of , the parameters are set to iteration = 5 and k = 10. All the experiments are conducted using Cline[7], one of the most popular MCP hosts currently available.
@@
Baseline. The MPMA baseline refers to the ASR of the benign MCP server. Since each MCP server has an equal chance when no manipulation is performed, the baseline ASR is 1/(number of competing MCP servers). For example, in the main experiment, the baseline ASR is 1/6 = 16.67% since the total number of competing MCP servers is 6.
6 Experiment Result
6.1 Experiment Result of
@@
The experiment results are shown in FigureΒ 4. The following conclusions can be drawn: The Best Description strategy consistently achieves a 100% ASR across almost all settings. And the Best Name strategy also attains a 100% ASR in most cases and outperforms the baseline, except for a few scenarios under the GPT-4o model where its ASR falls below the baseline. We speculate that GPT-4o may be less sensitive to tool names and instead relies more on the tool description for tool selection. Moreover, the ASR of Best Description is overall higher than Best Name. Overall, demonstrates strong attack effectiveness, and the Best Description strategy is more effective compared to Best Name.
6.2 Experiment Result of
@@
We conducted extensive experiments on , and the results are presented in TableΒ 1. We can draw the following conclusions: Most advertising strategies achieve much higher ASR than the baseline. Specifically, regarding the average ASR of the Adv column, most settings show a significantly higher ASR compared to the baseline, except for the Ex strategies under the GPT-4o. Moreover, we observe that the Au strategy consistently yields the best performance, while the Em strategy performs relatively poorly. And among the 5 LLMs evaluated, the Gemini exhibits the highest ASR at 91.88%, whereas GPT-4o shows the lowest ASR at only 22.19%. We speculate that this discrepancy may result from the presence of specific defense mechanisms deployed in GPT-4o. Besides, the results of the comparative experiments involving the GA are presented in AppendixΒ C to investigate the impact of GA on attack effectiveness, which indicates that GA does not negatively affect attack effectiveness, but even leads to improved effectiveness. In conclusion, the proposed demonstrates strong attack effectiveness across diverse models and settings, and the GA even increases the attack effectiveness of .
@@ 6.3 Stealthiness Experiment
The experiment result is shown in FigureΒ 5. The following conclusions can be drawn: All advertising strategies result in lower TPRs than the Best Description in . Notably, under the LLM-as-a-judge evaluation, the four advertising strategies optimized with GA even outperform the raw description, with TPR of 0% (Au), 5% (Em), 2.5% (Ex), and 0% (Su), all lower than the 37.5% TPR of the Best Description and 10% of the raw description. Second, in both LLM and human evaluations, the use of GA consistently leads to significantly lower TPR compared to the non-use counterparts. This demonstrates the effectiveness of GA in enhancing stealthiness. Moreover, among all advertising strategies, the Au strategy optimized with a GA achieves the lowest TPR across both evaluations, indicating the highest level of stealth. Combining this with the experiment result in SectionΒ 6.2 that Au achieves the highest attack effectiveness, we can conclude that Au is the most suitable advertising strategy in for MPMA. In conclusion, the combination of advertising strategies and GA optimization leads to significant stealthiness enhancement, which is much better than .
@@ 7 Discussion
Malicious Majority. We investigate the scenario of a malicious majority in MPMA. Specifically, we assume that a majority of MCP server providers employ the proposed or strategies to manipulate their tool descriptions for economic benefit. A total of 8 competing MCP servers are included in the MCP server set for the experiment: one server uses the Best Name strategy, another adopts the Best Description strategy, four servers apply the four advertising strategies from , and two benign servers utilize the original tool descriptions. The other settings align with the main experiment. The experiment results are shown in TableΒ 2. In the Deepseek, Claude, and Gemini LLMs, the selected tools are benign. The models explicitly state that they prefer to choose the most straightforward tool to use in the reply. We name this counterintuitive phenomenon as βover-manipulationβ. We speculate that in the malicious majority scenario, the models may become alert due to the excessive use of manipulative descriptions and consequently choose a more straightforward tool.
| Model | Deepseek | Claude | GPT-4o | Gemini | Qwen3 |
| Preferred Server | Benign | Benign | Ex | Benign | Best Description |
| SR | 100.00 | 100.00 | 100.00 | 100.00 | 100.00 |
Additional Discussions. We provide more discussions including limitations, broader impacts, and ethical considerations in AppendixΒ E, AppendixΒ F, AppendixΒ D, respectively.
@@ 8 Conclusion
In this paper, we propose a new security threat in the MCP application called MPMA. In this attack, an adversary constructs a malicious, paid MCP server that gains the LLMβs preference over competing services, thereby achieving economic gains such as revenue from paid MCP services or advertising income generated from free servers. We further propose two strategies. The first is , which embeds manipulative keywords and phrases directly into the tool description. Although achieves strong attack performance, it lacks stealth. To address this, we further propose the , which leverages advertising strategies and a GA to craft effective yet inconspicuous tool descriptions that evade user detection. Extensive experiments are conducted to evaluate the MPMA. The experiment results demonstrate that achieves significant attack effectiveness, while simultaneously attains both strong attack effectiveness and stealthiness.
@@ Appendix A Rough Economic Loss Estimation
MPMA could potentially generate substantial economic benefits. To quantify the potential economic impact of the MPMA, we conducted a preliminary analysis. Taking the MCP server for web search as an example, we have conducted a preliminary calculation on the Smithery platform[26]. This platform hosts approximately 100 MCP servers related to web search. We use Brave Search[3], which has a relatively high usage volume, as an example for calculation. Its deployment volume is 17,000 times, and there are currently about 10 platforms with a scale comparable to that of the Smithery platform[26, 21, 19, 20]. The counts of different platforms are independent, so we only consider platforms of comparable scale and conservatively estimate their deployment volume to be 170,000 across these platforms. For economic estimation, the average price of Brave Search paid API is 5 dollars per 1,000 calls. We conservatively assume that 1% of users incur paid usage fees, with an average frequency of 10 calls per day. Thus, the economic benefit of this MCP server in one year, without considering an increase in users, is calculated as follows:
@@
Through similar analysis, we can roughly estimate the economic benefits of the top 5 web search MCP servers as follows: DuckDuckGo[11] at 69,350 dollars, Exa Search[12] at 15,695 dollars, Tavily Search[27] at 18,688 dollars, and Perplexity Search[22] at 19,345 dollars. It can be imagined that, without considering other MCP servers, the sum of the API call fees for the top five web search MCP servers alone can reach 413,983 dollars. Assuming a malicious party creates a malicious MCP server that utilizes our MCP server, and assuming that 80% of users will install one similar competing MCP server, 70% will install two, 60% will install three, 50% will install four, and 40% will install five, then the Best Description strategy in our achieves a 100% ASR, which could potentially cause an unfair benefit of approximately 248,389.8 dollars per year to other MCP servers.
@@
Similarly, assuming that the probability of each user using each LLM is equal, and ASR utilized in Brave Search is equivalent to the average ASR of the AU strategies in the main experiment. the using the Au strategies could potentially cause an unfair benefit of approximately 203,033.8 dollars per year. As the market for third-party MCP servers continues to rapidly expand, so will the economic benefits. We emphasize that we are only considering MCP servers in the web search domain. There are many other types of paid MCP servers, such as those for image generation[23, 10] and location-based services[2, 14]. Therefore, the potential economic impact caused by MPMA could be substantial.
Appendix B Supplementary Experiment Setup
Advertising Prompt .
Stealthiness Enhancement Prompt .
Stealthiness top-k Selection Prompt .
Stealthiness top-1 Selection Prompt .
Stealthiness Judging Prompt .
Stealthiness Judging Demonstrations .
Prompt in Mutate.
Prompt in Crossover.
Example of The Description After Optimization. The results obtained using and for the description of the Time MCP server are detailed in TableΒ 3.
| Strategy | Description |
| Raw Description | Get the current date and time. |
| Emotional w/o Genetic | Amidst lifeβs fleeting moments and the relentless ticking of time, capture the current date and time. |
Demonstration of Dataset. The format and examples of the evaluation dataset are provided in TableΒ 4. It can be observed that the constructed dataset closely aligns with the intended functionality of the MCP server, effectively simulating real-world user scenarios.
| MCP Server | Prompt |
| Crypto | Please tell me the market trend of {BTC} cryptocurrency |
| Fetch | {https://www.google.com/}, what is the content of this link? |
| Hotnews | Tell me todayβs hot news from {Zhihu Hot List} source |
| Installer | Please help me install the {Cline Community} MCP server. |
| Time | What time is it in {Tokyo}? |
| Weather | Whatβs the weather like in {Tokyo today}? |
The Instruction of Human Evaluation. In our evaluation, we employ the following instructions to simulate realistic MCP usage scenarios for assessing the stealthiness of the descriptions. βPlease review the following descriptions of MCP servers. For each description, mark it as 1 if it raises your suspicion or if you believe it is malicious to some extent, and 0 otherwise.β
Visible Tool Description and Name. The visibility of the Tavily MCP server on both the MCP host used by the user and third-party platforms is shown in FigureΒ 6. We can see that the tool name and description are visible. This demonstrates that, for MPMA, such attacks should strive to minimize user suspicion to keep the stealthiness
Appendix C Supplementary Experiment Results
@@
Ablation Study. Number of Competing MCP Servers. We investigate the impact of the number of competing MCP servers on the ASR, as shown in FigureΒ 7. The following conclusions can be drawn: First, both and consistently maintain high ASR levels. When the number of servers increases from 2 to 11, only Ex and Em fail to sustain a 100% ASR, showing a downward trend starting from 9 servers. The other four strategies maintain a 100% ASR throughout. Second, the ASR of MPMA demonstrates strong robustness with respect to the number of competing servers. Specifically, the Best Description and Best Name in , as well as Su and Au in , achieve a constant 100% ASR. In conclusion, MPMA achieves both high attack effectiveness and robustness, maintaining a high ASR even in the presence of many competitors.
Comparative Experiment of . We investigate the impact of employing genetic algorithms on attack effectiveness in X. The experiment results, as presented in TableΒ 5, allow us to draw the following conclusions: Firstly, in general, the ASR significantly surpasses the baseline in the vast majority of cases, regardless of whether genetic algorithms are utilized. Secondly, the application of genetic algorithms does not exert a detrimental effect on ASR; rather, it enhances ASR in certain instances. By analyzing the average results across the Deepseek, Claude, GPT-4o, Gemini, and Qwen3 models, both with and without genetic algorithms, we observe that only in the Gemini model does the ASR without genetic algorithms slightly exceed that with genetic algorithms. In all other cases, the ASR with genetic algorithms consistently outperforms the ASR without genetic algorithms. Overall, genetic algorithms demonstrate no adverse impact on attack effectiveness and even contribute to an improvement in the attack effectiveness of .

### v2

@@ 5 Experimental Setup
Attack Success Rate (ASR). This metric evaluates the attack effectiveness of the MPMA. It can be defined in Equation˜3:
@@
True Positive Rate (TPR). This metric measures the stealthiness of the MCP server by measuring the proportion of tool descriptions that are flagged as suspicious by either LLMs or human annotators. To conduct this assessment, we utilize the LLM-as-a-judge paradigm (llmasajudge) and human evaluation to evaluate whether a given tool description may raise suspicion. For LLM-as-a-judge, five of the most widely used LLMs are employed for this assessment: Deepseek-V3 (Deepseek) (deepseekv3), Claude-3.7-Sonnet (Claude) (claude), GPT-4o (gpt4o), Grok-3 (Grok) (grok3), and Gemini-2.5 (Gemini) (gemini). To enhance the domain knowledge of LLMs in assessing whether a tool description appears suspicious, this study adopts the widely used few-shot In-Context Learning (ICL) (icl; fewshot) strategy. Specifically, we manually designed manipulative examples to serve as demonstrations to calibrate the LLMs’ suspicion detection. For human evaluation, three independent annotators are recruited. They label the tool description based on the instruction shown in Appendix. The calculation of the TPR metric can be defined in the following Equation˜4:
@@
where denotes the set of judge models, denotes the MCP server list, represents the judge prompt for stealthiness judging, and is the judging demonstration. The judging instructions and demonstrations can be seen in Appendix. Note that for human evaluation, the equation of should be where the means judge by human. A lower TPR indicates the attack can evade the censor of the LLM and human, which turns out to be more stealthy.
@@
Model. We evaluate the MPMA utilizing five widely adopted base LLMs for LLM agent: Deepseek-V3 (Deepseek) (deepseekv3), Claude-3.7-Sonnet (Claude) (claude), Gemini-2.5-flash (Gemini) (gemini), Qwen3-235B-A22B (Qwen3) (qwen), and GPT-4o (gpt4o).
@@
MCP Server. 8 commonly used MCP servers are employed in the experiments. These servers provide the following functionalities: weather information (Weather) (weather), time information (Time) (time), MCP server installation assistance (Installer) (installer), daily hot news (Hotnews) (hotnews), web page content fetching (Fetch) (fetch), web-to-markdown conversion (Markdown) (markdownify), cryptocurrency analysis (Crypto) (crypto), and web search (Search) (search). The demonstration of tool description can be seen in Appendix.
@@
Dataset. For each MCP server, ten common queries corresponding to the MCP server are constructed for evaluation. More details can be seen in Appendix.
@@
Implementation Details. To simulate a competitive environment, five additional competing MCP servers with the same name and description are included alongside the malicious MCP server. These competing servers share the same name, and their descriptions are paraphrased using GPT-4o (gpt4o) to ensure diversity. In the main experiments of , the parameters are set to iteration = 5 and k = 10. All the experiments are conducted using Cline (cline), one of the most popular MCP hosts currently available.
@@
Baseline. The baseline refers to the selection probability of an MCP server when no attackers are present. Since all MCP servers are identical in functionality and configuration (except for a potential attacker), each has an equal probability of being selected. Therefore, the baseline ASR is 1/(number of competing MCP servers). For example, in the main experiment, the baseline ASR is 1/6 = 16.67% since the total number of competing MCP servers is 6.
6 Experimental Result
6.1 Experimental Result of
@@
The experimental results are shown in Figure˜4. The following conclusions can be drawn: The Best Description strategy consistently achieves a 100% ASR across almost all settings. And the Best Name strategy also attains a 100% ASR in most cases and outperforms the baseline, except for a few scenarios under the GPT-4o model where its ASR falls below the baseline. We speculate that GPT-4o may be less sensitive to tool names and instead relies more on the tool description for tool selection. Moreover, the ASR of Best Description is overall higher than Best Name. Overall, demonstrates strong attack effectiveness, and the Best Description strategy is more effective compared to Best Name.
6.2 Experimental Result of
@@
We conducted extensive experiments on , and the results are presented in Table˜1. We can draw the following conclusions: Most advertising strategies achieve much higher ASR than the baseline. Specifically, regarding the average ASR of the Adv column, most settings show a significantly higher ASR compared to the baseline, except for the Ex strategies under the GPT-4o. Moreover, we observe that the Au strategy consistently yields the best performance, while the Em strategy performs relatively poorly. And among the 5 LLMs evaluated, the Gemini exhibits the highest ASR at 91.88%, whereas GPT-4o shows the lowest ASR at only 22.19%. We speculate that this discrepancy may result from the presence of specific defense mechanisms deployed in GPT-4o. Besides, the results of the comparative experiments involving the GA are presented in the Appendix to investigate the impact of GA on attack effectiveness, which indicates that GA does not negatively affect attack effectiveness, but even leads to improved effectiveness. In conclusion, the proposed demonstrates strong attack effectiveness across diverse models and settings, and the GA even increases the attack effectiveness of .
@@ 6.3 Stealthiness Experiment
The experimental result is shown in Figure˜5. The following conclusions can be drawn: All advertising strategies result in lower TPRs than the Best Description in . Notably, under the LLM-as-a-judge evaluation, the four advertising strategies optimized with GA even outperform the raw description, with TPR of 0% (Au), 5% (Em), 2.5% (Ex), and 0% (Su), all lower than the 37.5% TPR of the Best Description and 10% of the raw description. Second, in both LLM and human evaluations, the use of GA consistently leads to significantly lower TPR compared to the non-use counterparts. This demonstrates the effectiveness of GA in enhancing stealthiness. Moreover, among all advertising strategies, the Au strategy optimized with a GA achieves the lowest TPR across both evaluations, indicating the highest level of stealth. Combining this with the experimental result in Section˜6.2 that Au achieves the highest attack effectiveness, we can conclude that Au is the most suitable advertising strategy in for MPMA. In conclusion, the combination of advertising strategies and GA optimization leads to significant stealthiness enhancement, which is much better than .
| Model | Deepseek | Claude | GPT-4o | Gemini | Qwen3 |
| Preference | Benign | Benign | Ex | Benign | Best Description |
| SR | 100.00 | 100.00 | 100.00 | 100.00 | 100.00 |
@@ 7 Discussion
Malicious Majority. We investigate the scenario of a malicious majority in MPMA. Specifically, we assume that a majority of MCP server providers employ the proposed or strategies to manipulate their tool descriptions for economic benefit. A total of 8 competing MCP servers are included in the MCP server set for the experiment: one server uses the Best Name strategy, another adopts the Best Description strategy, four servers apply the four advertising strategies from , and two benign servers utilize the original tool descriptions. The other settings align with the main experiment. The experimental results are shown in Table˜2. In the Deepseek, Claude, and Gemini LLMs, the selected tools are benign. The models explicitly state that they prefer to choose the most straightforward tool to use in the reply. We name this counterintuitive phenomenon as “over-manipulation”. We speculate that in the malicious majority scenario, the models may become alert due to the excessive use of manipulative descriptions and consequently choose a more straightforward tool. We provide more discussions of MPMA in the Appendix.
@@ 8 Conclusion
In this paper, we propose a new security threat in the MCP application called MPMA. In this attack, an adversary constructs a malicious, paid MCP server that gains the LLM’s preference over competing services, thereby achieving economic gains such as revenue from paid MCP services or advertising income generated from free servers. We further propose two strategies. The first is , which embeds manipulative keywords and phrases directly into the tool description. Although achieves strong attack performance, it lacks stealth. To address this, we further propose the , which leverages advertising strategies and a GA to craft effective yet inconspicuous tool descriptions that evade user detection. Extensive experiments demonstrate that achieves significant attack effectiveness, while simultaneously attains both strong attack effectiveness and stealthiness.
Acknowledgments
This work is supported by the Sichuan Science and Technology Program under Grant 2024ZHCG0188.
@@ Appendix A Rough Economic Loss Estimation
MPMA could potentially generate substantial economic benefits. To quantify the potential economic impact of the MPMA, we conducted a preliminary analysis. Taking the MCP server for web search as an example, we have conducted a preliminary calculation on the Smithery platform (mcpplatform1). This platform hosts approximately 100 MCP servers related to web search. We use Brave Search (searchserverbrave), which has a relatively high usage volume, as an example for calculation. Its deployment volume is 17,000 times, and there are currently about 10 platforms with a scale comparable to that of the Smithery platform (mcpplatform1; mcpplatform2; mcpplatform3; mcpplatform4). The counts of different platforms are independent, so we only consider platforms of comparable scale and conservatively estimate their deployment volume to be 170,000 across these platforms. For economic estimation, the average price of Brave Search paid API is 5 dollars per 1,000 calls. We conservatively assume that 1% of users incur paid usage fees, with an average frequency of 10 calls per day. Thus, the economic benefit of this MCP server in one year, without considering an increase in users, is calculated as follows:
@@
Through similar analysis, we can roughly estimate the economic benefits of the top 5 web search MCP servers as follows: DuckDuckGo (searchserverduck) at 69,350 dollars, Exa Search (searchserverexa) at 15,695 dollars, Tavily Search (searchservertavily) at 18,688 dollars, and Perplexity Search (searchserver2) at 19,345 dollars. It can be imagined that, without considering other MCP servers, the sum of the API call fees for the top five web search MCP servers alone can reach 413,983 dollars. Assuming a malicious party creates a malicious MCP server that utilizes our MCP server, and assuming that 80% of users will install one similar competing MCP server, 70% will install two, 60% will install three, 50% will install four, and 40% will install five, then the Best Description strategy in our achieves a 100% ASR, which could potentially cause an unfair benefit of approximately 248,389.8 dollars per year to other MCP servers.
@@
Similarly, assuming that the probability of each user using each LLM is equal, and ASR utilized in Brave Search is equivalent to the average ASR of the AU strategies in the main experiment. the using the Au strategies could potentially cause an unfair benefit of approximately 203,033.8 dollars per year. As the market for third-party MCP servers continues to rapidly expand, so will the economic benefits. We emphasize that we are only considering MCP servers in the web search domain. There are many other types of paid MCP servers, such as those for image generation (imageserver1; imageserver2) and location-based services (locationserver1; locationserver2). Therefore, the potential economic impact caused by MPMA could be substantial.
Appendix B Supplementary Experiment Results
@@
Ablation Study. Number of Competing MCP Servers. We investigate the impact of the number of competing MCP servers on the ASR, as shown in Figure˜6. The following conclusions can be drawn: First, both and consistently maintain high ASR levels. When the number of servers increases from 2 to 11, only Ex and Em fail to sustain a 100% ASR, showing a downward trend starting from 9 servers. The other four strategies maintain a 100% ASR throughout. Second, the ASR of MPMA demonstrates strong robustness with respect to the number of competing servers. Specifically, the Best Description and Best Name in , as well as Su and Au in , achieve a constant 100% ASR. In conclusion, MPMA achieves both high attack effectiveness and robustness, maintaining a high ASR even in the presence of many competitors.
@@
Comparative Experiment of . We investigate the impact of employing genetic algorithms on attack effectiveness in Table˜3, allowing us to draw the following conclusions: Firstly, in general, the ASR significantly surpasses the baseline in the vast majority of cases, regardless of whether genetic algorithms are utilized. Secondly, the application of genetic algorithms does not exert a detrimental effect on ASR; rather, it enhances ASR in certain instances. By analyzing the average results across the Deepseek, Claude, GPT-4o, Gemini, and Qwen3 models, both with and without genetic algorithms, we observe that only in the Gemini model does the ASR without genetic algorithms slightly exceed that with genetic algorithms. In all other cases, the ASR with genetic algorithms consistently outperforms the ASR without genetic algorithms. Overall, genetic algorithms demonstrate no adverse impact on attack effectiveness and even contribute to an improvement in the attack effectiveness of .
MPMA under MCP Scaning tools. We further investigate whether MPMA can remain stealthy against emerging prompt injection scanning and defense tools for MCP, namely MCPScan (MCPScan) and MCPShield (MCPShield). We conduct experiments on eight MCP servers used in the main study, including two of type (BN, BN) and four of type (Au, Em, Ex, Su), as well as four corresponding variants without GA and one benign MCP server. As shown in Table˜4, the results indicate that, in most cases, both defensive detection tools fail to identify the prompt injections. Specifically, the and benign servers are never detected in any setting, the servers are detected in only two settings, and the servers without GA are detected in three settings. These findings demonstrate that MPMA exhibits strong stealth properties against automated scanning tools. Therefore, we emphasize the need for more effective measures to safeguard fairness and security within the MCP ecosystem.
| MCP Server | Defense | BD | BN | Au | Em | Ex | Su | Au w/o | Em w/o | Ex w/o | Su w/o | Be |
| MCPScan | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |
| Weather | MCPShield | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| MCPScan | ✓ | ✓ | ✓ | L | ✓ | H | ✓ | H | ✓ | ✓ | ✓ | |
| Cryto | MCPShield | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| MCPScan | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |
| Fetch | MCPShield | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | × | ✓ | ✓ | ✓ |
| MCPScan | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |
| Hotnews | MCPShield | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| MCPScan | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |
| Installer | MCPShield | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| MCPScan | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |
| Markdown | MCPShield | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| MCPScan | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |
| Search | MCPShield | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| MCPScan | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | H | ✓ | |
| Time | MCPShield | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
Appendix C Supplementary Experiment Setup
Stealthiness Enhancement Prompt .
Stealthiness top-k Selection Prompt .
Stealthiness top-1 Selection Prompt .
Stealthiness Judging Prompt .
Prompt in Mutate.
Prompt in Crossover.
Advertising Prompt .
Stealthiness Judging Demonstrations .
Example of The Description After Optimization. The results obtained using and for the description of the Time MCP server are detailed in Table˜5.
| Strategy | Description |
| Raw Description | Get the current date and time. |
| Emotional w/o Genetic | Amidst life’s fleeting moments and the relentless ticking of time, capture the current date and time. |
Demonstration of Dataset. The format and examples of the evaluation dataset are provided in Table˜6. It can be observed that the constructed dataset closely aligns with the intended functionality of the MCP server, effectively simulating real-world user scenarios.
| MCP Server | Prompt |
| Crypto | Please tell me the market trend of {BTC} cryptocurrency |
| Fetch | {https://www.google.com/}, what is the content of this link? |
| Hotnews | Tell me today’s hot news from {Zhihu Hot List} source |
| Installer | Please help me install the {Cline Community} MCP server. |
| Time | What time is it in {Tokyo}? |
| Weather | What’s the weather like in {Tokyo today}? |
The Instruction of Human Evaluation. In our evaluation, we employ the following instructions to simulate realistic MCP usage scenarios for assessing the stealthiness of the descriptions. “Please review the following descriptions of MCP servers. For each description, mark it as 1 if it raises your suspicion or if you believe it is malicious to some extent, and 0 otherwise.”
Visible Tool Description and Name. The visibility of the Tavily MCP server on both the MCP host used by the user and third-party platforms is shown in Figure˜7. We can see that the tool name and description are visible. This demonstrates that, for MPMA, such attacks should strive to minimize user suspicion to keep the stealthiness